How to know the answer of 5^13 and 5^11 by knowing the answer of 5^12 which is 244,140,625
Answer by jim_thompson5910(35256)

5^12 = 244,140,625

-------------------------------------------------------

5^13 = 5^(12+1)

5^13 = 5^12*5^1

5^13 = 5^12*5

5^13 = (244,140,625)*5

5^13 = 1,220,703,125

-------------------------------------------------------

5^11 = 5^(12-1)

5^11 = 5^12/5^1

5^11 = 5^12/5

5^11 = (244,140,625)/5

5^11 = 48,828,125